Chapter 5, Section 1, Exercise 012 Incorrect. The following is a set of hypotheses, some information from one or more samples, and a standard error from a randomization distribution. Test Ho : ?? = ?2 vs Ha : ? > ?2 when the samples have n = n, = 40, ??35.7, x, = 33.1, s? = 1.24, and $2 1.11. The standard error of Ti from the randomization distribution is 0.26. Find the value of the standardized z-test statistic. Round your answer to two decimal places. 8.75 the absolute tolerance is +/-0.01
Answer #1

The Standardized Test Statistic Z = (\bar{x}_1 - \bar{x}_2)/Standard Error

Z = (35.7 - 33.1)/0.26 = 2.6/0.26 = 10.00